What volume of a 25% (w/v) solution of HNO3?
1
Expert's answer
2020-02-06T06:27:09-0500

Solution.

"w(HNO3) = \frac{m(HNO3)}{m(solution)} \times 100 \%"

"m(solution) = \frac{m(HNO3) \times 100 \%}{w(HNO3)}"

m(solution) = 50.4 g

"m(H2O) = m(solution) - m(HNO3)"

m(H2O) = 37.8 g

Answer:

m(solution) = 50.4 g